Find the vertex of the graph of the function.

f(x) = (x + 4)2 - 1

(0, -4)

( -1, 0)

( -4, -1)

( -1, -4)

show work
1
Expert's answer
2014-06-10T03:13:05-0400
f(x) = (x + 4)2- 1=x2+8x+15
The vertex can be found using the formula x0 = -b/2a = -8/2 = -4;
y0 = f(-4) = -1
Answer: (-4;-1)